OpenStudy (anonymous):

x cos y + y cos x = 1 at (0,1) implicit differentiation

OpenStudy (ipwnbunnies):

dy/dx at (0,1)?

OpenStudy (kainui):

Seems reasonable. Show your best attempt peper! I'll show you where you need help!

OpenStudy (anonymous):

An equation of the line tangent to the graph of x cosy + y cosx = 1 at the point (0,1) So I think dy/dx at (0,1)

OpenStudy (ipwnbunnies):

Yeah, show us your best attempt. :)

OpenStudy (ipwnbunnies):

Me thinks you'll have to use Product Rule twice.

OpenStudy (anonymous):

all I have is: x(-siny dy/dx) + (cosy)(1) + y (-sinx) + (cosx)(dy/dx)

OpenStudy (kainui):

I like to think of implicit differentiation as really just normal differentiation you've done all along. For instance, let's look at: y=x Obviously, we know that y'=1 here. But a closer look we notice that: \[y^1=x^1\]\[\frac{ d }{ dx }(y^1)=\frac{ d }{ dx }(x^1)\]\[1*y^{1-1}*\frac{ dy }{ dx }=1*x^{1-1}*\frac{ dx }{ dx }\]notice how I just did the normal power rule where I take the power, multiply it out front, subtract 1 from the exponent? Then I multiplied by the derivative of the inside function with the chain rule. BUT I did exactly the same thing to both sides. However now we can clean this up a bit! What is the 1-1 power on both y and x? 0! And anything to the zero power is just 1. So we can just leave ourselves with: \[\frac{ dy }{ dx }=\frac{ dx }{ dx }\] And now, what about that dx/dx term? Well, how much does x change when you change x? Obviously there's a 1 to 1 correspondence going on here! So that too is 1. \[\frac{ dy }{ dx }=1\] --- So now try this with something else, like \[y=x^2\] and if you reorganize it to look like this: \[y^{1/2}=x\] they will both have the same derivative!

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

A. (cos1)x+y=1 B. x+y=1 C.-(sin1)x+y=1 D.x-y=1 E. (tan1)x+y=1

OpenStudy (ipwnbunnies):

Err, if you go back to your expanded derivative, you'll have to set it equal to 0, since you took the derivative of both sides of your original equation. Then, try to get the terms that do not have (dy/dx) on the right side. Then, factor out the (dy/dx) and bring the remains to the right side. Now, you'll have an equation for dy/dx. Plug in the point.

OpenStudy (anonymous):

I tried.. Twice in fact! And somehow managed to end up with 0 in the denominator.

OpenStudy (ipwnbunnies):

Hmm, doing it quick in my head, Idk how you're getting 0 in the denominator if the point is (0,1).

OpenStudy (ipwnbunnies):

Write out the equation for dy/dx again. I'm definitely not getting 0 in denominator.

OpenStudy (anonymous):

Okay. So I do the product rule twice and get x(-siny dy/dx) + (cosy)(1) + y (-sinx) + (cosx)(dy/dx) = 0 so then I move things around and end up with: dy/dx (-sinxy + cosx) = -coxy + sinxy have I messed up in these steps?

OpenStudy (ipwnbunnies):

On the left, you should have dy/dx (-siny*x + cosx)* ;D Maybe that's where you messed up

OpenStudy (ipwnbunnies):

You wrote y*-sinx, instead of x*-siny if you didn't see yet

OpenStudy (anonymous):

what you said I should have on the left I do have.. don't I? I just failed to put the *

OpenStudy (anonymous):

or where are you getting the y*-sinx instead of x*-siny

OpenStudy (ipwnbunnies):

Look at what you wrote when you factored out the dy/dx.

OpenStudy (ipwnbunnies):

(-sinyx = cosx) should be inside the parentheses

OpenStudy (ipwnbunnies):

(-sinyx + cosx)*

OpenStudy (anonymous):

on the left side?

OpenStudy (ipwnbunnies):

Yeah

OpenStudy (ipwnbunnies):

Now get dy/dx alone. You shouldn't get 0 in denominator now. You flipped the variables in one term, small mistake.

OpenStudy (anonymous):

dy/dx (-sinxy + cosx) is what I had and you had (-sinyx + cosx) can you explain to me how I am suppose to plug this in. I think I am misunderstanding, I had dy/dx = (-cosy +sinxy)/(-sinyx + cosx) dy/dx = (-cos(1) + sin(0)(1)) / (-sin(1)(0) + cos(0)) dy/dx = (-cos1) / 0 soooooooo I have know idea, I guess, how or what I am supposed to do because I am obviously doing something incorrect.

OpenStudy (ipwnbunnies):

Err, there's no 0 in the denominator silly! Look at the cosine again.

OpenStudy (ipwnbunnies):

You get 0 for the first term, plus the cosine of 0...:OOO

OpenStudy (anonymous):

.... and the cosine of 0 is 1?

OpenStudy (ipwnbunnies):

Yes, yes it is. Congrats, that was almost painful.

OpenStudy (anonymous):

A. (cos1)x+y=1 B. x+y=1 C.-(sin1)x+y=1 D.x-y=1 E. (tan1)x+y=1 sooooo my answer doesn't look like any of the optons... so thats painful haha

OpenStudy (anonymous):

i like pain but i dont like implicit diff

OpenStudy (anonymous):

options*

OpenStudy (ipwnbunnies):

Implicit differentiation isn't hard when you know the Implicit Function Theorem ;)

OpenStudy (anonymous):

it's not hard, i just heavily dislike it

OpenStudy (ipwnbunnies):

Whoa, what. Those choices don't look right. This probably is gonna be painful. I didn't do the derivative yet.

OpenStudy (ipwnbunnies):

I don't have time to go over it again :(

OpenStudy (ipwnbunnies):

OHOHHHOH, you need to find a tangent line to the point.

OpenStudy (anonymous):

I showed you the choices earlier! And that's alright, my exam will just kill me tomorrow.

OpenStudy (ipwnbunnies):

Well, we just found the slope of the tangent line at the point. Use point-slope equation now. :D

OpenStudy (ipwnbunnies):

What'd you get for dy/dx, first?

OpenStudy (ipwnbunnies):

Oh...We did the hard part. We just gotta find the equation of the line. Algebra 1 stuff.

OpenStudy (ipwnbunnies):

y - 1 = -cos(1)(x - 0) y = -xcos(1) + 1

OpenStudy (ipwnbunnies):

y + xcos(1) = 1, A. Good night.

OpenStudy (anonymous):

You are the best! Thank you so much for your help!

OpenStudy (ipwnbunnies):

It'd be nice to say 'no prob,' but that was truly painful.

OpenStudy (ipwnbunnies):

Good luck on your test.

OpenStudy (anonymous):

That it was. My professor insists we do Purdue University problems because they are challenging. Thank you so much again! Hopefully it will be smooth sailing.

OpenStudy (ipwnbunnies):

Ahh yes, Purdue has a good math/engineering department. Cheers.